# Structured Kolmogorov-Arnold Neural ODEs for Interpretable Learning and Symbolic Discovery of Nonlinear Dynamics

### v2

@@ 5 Conclusion
Comprehensive experiments on multiple nonlinear dynamical systems demonstrate that SKANODE outperforms existing NODE variants in both predictive accuracy and interpretability. These results highlight the importance of combining inductive structural biases with symbolic discovery capabilities to guide the learning process toward solutions that are physically consistent, robust, and interpretable. Specifically, SKANODE accurately recovers governing equations that align with ground truth on classical nonlinear benchmarks, including the Duffing and Van der Pol oscillators. In the real-world F-16 aircraft system, SKANODE identifies complex hysteretic behavior and derives an interpretable symbolic equation that captures this nonlinear phenomenon.
The ability to uncover governing equations from indirect observations makes SKANODE particularly well-suited for scientific and engineering domains where full-state measurements are often impractical, and model transparency is critical for trustworthy decision-making. Future work will explore the integration of domain-specific physics priors and constraints within the symbolic extraction process.
Another important direction concerns robustness to real-world sensing imperfections such as sensor drift, measurement bias, and colored noise. While SKANODE has been validated under varying white noise levels on simulated systems such as the Duffing oscillator, future work may further investigate SKANODE’s ability to extract latent dynamics from measurements affected by structured or colored noise. One promising extension is to incorporate a decoupled observation model, where the output mapping is trained with the dynamics to account for sensing distortions. Such modeling flexibility would allow SKANODE to adapt to more challenging real-world sensing environments.

### v3

@@ 5 Conclusion
Comprehensive experiments on multiple nonlinear dynamical systems demonstrate that SKANODE outperforms existing NODE variants as well as classical system identification baselines in both predictive accuracy and interpretability. Specifically, SKANODE recovers equation-level descriptions that match the expected nonlinear mechanisms across all considered systems. For the Duffing oscillator, the discovered symbolic form captures the characteristic cubic stiffness nonlinearity, while the structured latent coordinates evolve as physically meaningful displacement and velocity. For the Van der Pol oscillator, the extracted symbolic structure reflects the nonlinear damping mechanism through the expected state-dependent coupling. For the real-world F-16 ground vibration dataset, the inferred latent phase portraits exhibit closed-loop hysteretic patterns, and the learned symbolic model captures an associated internal hysteretic mechanism via the additional latent hysteretic state. In all three cases, the proposed method yields more accurate and robust predictions than black-box NODE baselines and classical ARX/NARX identification. These results highlight the importance of combining inductive structural biases with symbolic discovery capabilities to guide the learning process toward solutions that are physically consistent, robust, and interpretable.
The ability to uncover governing equations from indirect observations makes SKANODE particularly well-suited for scientific and engineering domains where full-state measurements are often impractical, and model transparency is critical for trustworthy decision-making. In terms of applicability, the proposed framework is intended for nonlinear dynamical systems whose dominant behavior can be reasonably captured by a smooth ODE that admits a structured state-space representation (e.g., second-order dynamics with displacement and velocity states, or modest extensions with additional latent internal variables). This class includes a wide range of engineering systems in structural dynamics, vibration mitigation, aeroelasticity, and electromechanical devices. Effective use further relies on a measurement model compatible with the imposed structure (e.g., acceleration measurements consistent with second-order dynamics), on sufficient observability of the structured latent states from the available measurements, and on sufficiently informative excitation and, where relevant, known external inputs. When the above conditions are violated—for example in systems dominated by non-smooth switching or impacts, strong sensing distortions such as bias, drift, or colored noise, unmeasured forcing, or weak excitation—the recovered latent states and symbolic expressions may lose physical meaning or become non-unique. In such cases, additional domain-specific constraints, hybrid formulations, or more flexible observation models may be required.
Future work will explore the integration of stronger domain-specific priors and constraints within the symbolic extraction process, as well as extensions that improve robustness to realistic sensing imperfections. One promising extension is to incorporate a partially decoupled observation model trained jointly with the dynamics to account for sensing distortions. Such modeling flexibility would allow the framework to adapt to more challenging real-world sensing environments.